two angles are supplementary. the measure of the smaller angle is one-half the measure of the larger angle. Find the degree measure of the larger angle.
Answer by oberobic(2304)

x + y = 180
so
x = 180-y
.
x = 2y
.
x=x
so
.
2y = 180-y
3y = 180
y = 60
.
x = 120
.
Done.